# Key-value memory in the brain

### v1

@@ 1 Introduction
Despite the apparent fragility of memory, there is no decisive evidence that information, once stored, is ever permanently lost. The storage capacity of the brain of course must be finite, but that does not seem to be the principal limiting factor on memory performance. Rather, it is the retrieval process that fundamentally limits performance: the relevant information may be there, but cannot always be found (Tulving,, 1974; Crowder,, 1976; Lewis,, 1979; Miller,, 2021). Some of the evidence supporting this view will be summarized below.
A retrieval-oriented view of memory performance places the primary explanatory burden on how memories are addressed (i.e., how the retrieval system keeps track of storage locations), and how they are queried (i.e., how the retrieval system maps sensory cues to addresses). There is a long history of theorizing about these concepts in cognitive psychology and neuroscience (Kahana,, 2012; Chaudhuri and Fiete,, 2016). Recently, the field of machine learning has developed its own analysis of these concepts, which form the basis of high-performing systems like transformers and fast weight programmers (Vaswani et al.,, 2017; Schmidhuber,, 1992). There is increasing recognition that a significant aspect of intelligence (in both natural and artificial systems) is effective information retrieval (Graves et al.,, 2016; Gershman and Daw,, 2017; Geva et al.,, 2021; Irie et al.,, 2022; Allen-Zhu and Li,, 2024).
@@ 2.1 From correlations to kernels
One of the earliest formalizations of a key-value memory was Kohonen’s correlation matrix memory model (Kohonen,, 1972), which was subsequently used by Pike, (1984) to explain a range of human memory phenomena. Here we slightly change the notation and terminology to bring this model into correspondence with more recent formalizations (Figure 1, left). Each input, indexed by $n$, consists of a key vector $\mathbf{k}_{n}$ and a value vector $\mathbf{v}_{n}$ (which we take to be row vectors). Intuitively, the keys encode information about memory addresses (hence we will refer to the set of possible keys as the address space), while the values encode information about memory content. These two representations are linked in memory by an “associator” matrix $\mathbf{M}$, which is initialized at 0 and incremented by the outer product of the key and value vectors after each input is presented:
@@
It is easy to see that this is just simple Hebbian learning between key and value units encoding the vector elements. In neurobiology, the standard interpretation is that $M_{ij}$ is encoded by the synaptic strength between neurons representing value element $j$ and key element $i$. The synaptic strength is increased when the two neurons fire coincidentally (Caporale and Dan,, 2008). Needless to say, learning in the brain is more complicated than this, but for present purposes we will take for granted the biological plausibility of Eq. 1 (see also Limbacher and Legenstein,, 2020, for a related, but more complex, learning rule).
@@
The correlation matrix memory is heteroassociative because $\mathbf{M}$ stores information about the relationship between two different kinds of objects or object properties (see also Steinbuch and Piske,, 1963, for one of the earliest such models). If we impose the constraint that keys and values are the same, we get an autoassociative memory (Willshaw et al.,, 1969; Anderson,, 1970; Amari,, 1972; Nakano,, 1972). This idea is now most closely associated with the work of Hopfield, (1982), and such models (when the inputs are binary) are typically referred to as Hopfield networks.
@@
In neural network terms, this is equivalent to activating pattern $\mathbf{q}$ on the key units, which produces an activity pattern in downstream value units based on the learned synaptic strengths ($\mathbf{M}$). The retrieved value vector $\hat{\mathbf{v}}$ combines stored values associated with keys that are similar to the query vector. To see this, it is helpful to rewrite Eq. 2 in its “dual” form (Ba et al.,, 2016; Katharopoulos et al.,, 2020):
@@
where $\{\alpha_{n}\}$ is a set of “attention weights” computed by $\alpha=\sigma(S(\mathbf{K},\mathbf{q}))=\mathbf{q}\mathbf{K}^{\top}$, and $\mathbf{K}$ is the matrix consisting of all $N$ key vectors (i.e., its rows are $\mathbf{k}_{n}$). The function $S(\cdot,\cdot)$ is a similarity kernel (linear in this case), expressing the match between keys and values, and $\sigma(\cdot)$ is a separation operator (the identity function in this case), pushing apart similar memories. The retrieved value vector is thus a weighted combination of stored values, where the attention weights correspond to the similarity between the query and each key, mapped through a separation operator.
The dual form is useful because it shows how the model can be straightforwardly generalized by considering other similarity kernels and separation operators. Any positive semidefinite kernel can be expressed as the inner product in some vector space (Schölkopf,, 2002), which means that alternative similarity functions can be constructed by mapping the keys and queries through a feature transform $\phi(\cdot)$ applied to each row:
@@
It is also possible to construct kernels with infinite-dimensional vector spaces by working directly with an explicitly defined similarity kernel, such as the widely used radial basis function kernel, which Tsai et al., (2019) found to produce the best results.
The generalized correlation matrix memory can implement the celebrated self-attention mechanism in transformers (Vaswani et al.,, 2017):
@@
where $D$ is the dimensionality of the key/value vectors, and the softmax function is used as the separation operator. Here we have used $\tilde{\alpha}$ to denote the attention weights prior to the application of the separation operator. When $\sigma(\cdot)$ is set to the identity function, we recover linearized attention (Katharopoulos et al.,, 2020), which can be expressed as a special form of recurrent neural network known as a fast weight programmer (Schmidhuber,, 1992; Schlag et al.,, 2021). Linearized attention has a potential computational advantage over softmax attention, because its recurrent form allows for linear-complexity sequence processing, while giving up the parallel computation property of the attention form (whose complexity, however, is quadratic in the number of sequence elements). Choromanski et al., (2021) and Peng et al., (2021) showed how to rigorously approximate softmax attention using random features.
As shown by Millidge et al., (2022), a variety of classical memory models can also be derived by different choices of similarity kernel and separation function. For example, sparse distributed memory (Kanerva,, 1988) can be obtained by setting $\sigma(\cdot)$ to a threshold function, and dense associative memory (Krotov and Hopfield,, 2016) can be obtained by setting $\sigma(\cdot)$ to a rectified polynomial.
In the noiseless setting, the ideal separation function is the max operator, since this will always return the stored value associated with the matching key (Millidge et al.,, 2022). However, this separation function is not robust to noise—small perturbations can cause large retrieval errors. Thus, the design of a memory system needs to balance separability and robustness. In addition, the memory may be used for generalization to new inputs (e.g., Vaswani et al.,, 2017; Ramsauer et al.,, 2021; Bricken et al.,, 2023), in which case perfect matching is not the goal. For example, key-value memories have been used extensively for question-answering tasks (Sukhbaatar et al.,, 2015; Miller et al.,, 2016), where $\mathbf{q}$ represents a question and $\mathbf{v}$ represents an answer. The system stores a set of question-answer pairs and tries to use this knowledge base to answer novel questions.
@@ 2.2 Representational structure
In some systems, the key and query mappings are assumed to be fixed, either randomly or based on some regular structure. In these systems, the address space is conceived as a kind of “scaffold” for the indexing of information content. In the sparse distributed memory model (Kanerva,, 1988), the scaffold is random, so that similar keys do not systematically index similar values. This mimics (approximately) the organization of random access memory in digital computers. A familiar example of a structured scaffold is the alphabetical index found at the end of books. In the Hopfield network (Hopfield,, 1982) and related autoassociative memory models, the scaffold is identical to the value space. This property endows the models with content addressability (Kohonen,, 1980): memory addresses are accessed by matching values directly to queries. Most psychological and neural models of memory share this property, though they implement it in different ways.
Some models assume that the scaffold is randomly constructed in such a way that it carries a structural imprint. For example, if inputs are assigned to random addresses, but these addresses change slowly over time, then inputs experienced nearby in time will encoded as more “similar” in the address space. Temporal autocorrelation, in the absence of additional structure, is able to account for many aspects of human (Landauer,, 1975) and animal (Estes,, 1955) memory.
@@ 2.3 The ubiquity of key-value memory
The work reviewed above focused on explicit constructions of key-value memory. It turns out that other models are sometimes implicitly equivalent. Irie et al., (2022) showed that linear layers trained by gradient descent (ubiquitous in many machine learning models) can also be expressed as key-value memories. Here we briefly review this theorem.
@@
where $\mathbf{W}_{0}$ is the initial weight matrix, and $\mathbf{e}_{n}=-\eta_{n}(\nabla_{\mathbf{y}}\mathcal{L})_{n}$ is the error signal with learning rate $\eta_{n}$ and loss function $\mathcal{L}$. Generalizing a classic result on kernel machines (Aizerman et al.,, 1964), Irie et al., (2022) showed that this construction is equivalent (in the sense of producing the same output for a given input) to the following linear key-value memory:
@@
where (using our earlier notation) $\alpha=\mathbf{q}\mathbf{K}^{\top}$ with $\mathbf{v}_{n}=\mathbf{e}_{n}$, $\mathbf{k}_{n}=\mathbf{x}_{n}$, and $\mathbf{q}=\mathbf{x}$, given an arbitrary input $\mathbf{x}$. Thus, linear layers effectively memorize their experienced error patterns, computing their outputs as linear functions of these memories. This interpretation is intriguing in light of the fact that errors are particularly salient in human memory (Green,, 1956; Hirshman et al.,, 1989; Sakamoto and Love,, 2004; Rouhani et al.,, 2018; Bein et al.,, 2021).
The linear layers retain information about all training inputs—they never “forget” (Irie et al.,, 2022). However, retrieval access may be (transiently) lost. We discuss evidence for this idea from psychology and neuroscience below.
@@ 3 Neurobiological substrates
While key-value memories are loosely inspired by the brain, it remains an open question how to implement them in a biologically plausible manner. Eq. 1 is a (somewhat) biologically plausible rule for learning associations between keys and values, but we still need rules for storing the keys and values themselves. In the section on representation learning, we described the widely used approach of modeling keys, queries, and values as linear mappings from input vectors (which themselves could be learned). This implies additional learning rules for the linear mappings. One could posit that they are learned by backpropagating the errors from whatever objective function is being optimized for a downstream task (see, for example, Limbacher and Legenstein,, 2020). This would require a biologically plausible approximation of the backpropagation algorithm (Lillicrap et al.,, 2020).
Alternatively, Tyulmankov et al., (2021) have proposed a non-Hebbian learning rule for key learning. They view the key-value memory as a three-layer neural network, where the input $\mathbf{x}$ (first layer) is transformed into a pattern of attention $\boldsymbol{\alpha}$ (hidden layer), which is finally transformed into retrieved values $\hat{\mathbf{v}}$ (output layer); see the right panel of Figure 1. Each hidden layer unit represents a “slot” to which a single input (or possibly multiple inputs) gets assigned. In this view, the key matrix $\mathbf{K}$ corresponds to the input-hidden synaptic strengths, and the value matrix $\mathbf{V}$ (where row $n$ corresponds to value vector $\mathbf{v}_{n}$) corresponds to the hidden-output synaptic strengths. The proposed learning rule for the synapse connecting input unit $j$ to hidden unit $i$ is (simplifying slightly, and dropping time indices):
@@
where $\mu\in\{0,1\}$ is a global third factor (possibly a neuromodulator like acetylcholine or dopamine) and $\gamma_{i}\in\{0,1\}$ is a local third factor (possibly a dendritic spike). The factors are binary to promote sparsity in the hidden layer. The local third factor tags the least-recently-used hidden unit as eligible for plasticity. Tyulmankov et al., (2021) point out that the key learning rule resembles behavioral time scale plasticity (Bittner et al.,, 2017), which has been observed in the hippocampus. The possibility that key learning occurs in the hippocampus will be considered in detail below.
For value learning, Tyulmankov et al., (2021) clamp the output layer to the target value $\mathbf{v}$ and then update the synapse connecting hidden unit $i$ to output unit $m$ according to:
@@
Kozachkov et al., (2023) have proposed a different kind of architecture based on the “tripartite synapse” consisting of pre-synaptic and post-synaptic neurons modulated by an astrocyte (a type of glial cell). In particular, they posit that this motif applies to the hidden-to-output synapses in the three-layer network implementation of key-value memory. The activation of each astrocyte process is modeled as a linear function of the hidden unit activations. Kozachkov et al., (2023) show how the astrocyte processes collectively compute the similarity function $S(\mathbf{K},\mathbf{q})$, which then multiplicatively modulate the hidden-to-output weights so that the network as a whole implements the transformer self-attention described earlier.
Box 1: Random projections of attractor states for error-correcting key-query matching
@@
Vector-HaSH (Chandra et al.,, 2023) and MESH (Sharma et al.,, 2022) are tripartite networks consisting of an input (sensory) layer, a densely connected layer, and a layer consisting of a modular set of fixed recurrent attractor networks. These models can be directly connected with a key-query-value formulation: Define the tripartite network weights to be ${\bf W}_{sd},{\bf W}_{da},{\bf W}_{ad},{\bf W}_{ds}$ (from the sensory layer $s$ to a densely connected layer $d$, dense to attractor layer $a$, attractor to dense, and dense to sensory, respectively). Define the attractor, dense layer, and attractor states to be $\{{\bf a}\},\{{\bf d}\},\{{\bf s}\}$, respectively. If we view the dense layer responses as linear (in practice they are nonlinear), the network performs the following operation: $\hat{{\bf s}}=\phi({\bf s}{\bf W}_{sd}{\bf W}_{da},\{{\bf a}\}){\bf W}_{ad}{\bf W}_{ds}$ where $\phi(.,.)$ is an efficient nearest-neighbor (NN) operation performed on the sensory inputs by the scaffold consisting of the recurrent attractor network and the dense layer. The attractor retrieves the nearest of the attractor states $\{{\bf a}\}$ to the input ${\bf s}{\bf W}_{sd}{\bf W}_{da}$. This model can be interpreted as the key-value system of Figure 1B, in which the output of the attractor layer can be defined as a (denoised) query ${\bf q}_{n}={\bf x}_{n}=\phi({\bf s}{\bf W}_{sd}{\bf W}_{da},\{{\bf a}\})$, and ${\bf W}_{ad}$ and ${\bf W}_{ds}$ correspond to the key and value matrices, respectively (i.e., $\mathbf{K}={\bf W}_{ad}^{\top}$ and $\mathbf{V}={\bf W}_{ds}$); that is, the diagram’s hidden states (attention weights) are $\alpha_{n}={\bf x}_{n}{\bf W}_{ad}$. The combined attractor-dense layer scaffold computes a similarity $\sigma(S({\bf K},{\bf q}))$ and the network produces the output $\hat{{{\bf v}}}_{n}=\hat{{\bf s}}_{n}=\alpha_{n}{\bf W}_{ds}={\bf x}_{n}{\bf W}_{ad}{\bf W}_{ds}=\phi({\bf s}_{n}{\bf W}_{sd}{\bf W}_{da},\{{\bf a}\}){\bf W}_{ad}{\bf W}_{ds}=\sigma(S({\bf K},{\bf q}_{n})){\bf W}_{ds}$. The value weights (from the dense to the sensory layer) and part of the query weights (from the sensory to the dense layer) are trained by a Hebbian rule, so that ${\bf W}_{sd}=\sum_{n}{\bf s}_{n}^{\top}{\bf d}_{n}$ and ${\bf W}_{ds}=\sum_{n}{\bf d}_{n}^{\top}{\bf s}_{n}$; the weights ${\bf W}_{da}$ are random and fixed, and ${\bf W}_{ad}$ are also held fixed (after being set as a pseudoinverse on the random ${\bf W}_{da}$, so these may also be viewed as random). Thus, the key weights are random and fixed; the choice of modular attractor states and random projections from attractor neurons to the dense layer pre-defines keys as a well-separated set so that NN behaves well for noisy inputs. The query weights are partly random, fixed, and coupled with the keys, and partly learned through Hebbian plasticity.
@@
Recently, Sharma et al., (2022) and Chandra et al., (2023) presented neurobiologically motivated models (MESH and Vector-HaSH) that use random projections of modular fixed point (attractor) networks to address memories (see Box 3 for details). The attractor modules are fixed and not learned. The fixed points within each module may be sparse and random (Sharma et al.,, 2022) or possess a 2-dimensional geometry (Chandra et al.,, 2023), with hexagonally periodic activity patterns like grid cells, which are found in the medial entorhinal cortex (Hafting et al.,, 2005). These modular attractor networks randomly and densely project into a larger network (which in Vector-HaSH corresponds to the hippocampus) with self-consistent return projections, forming a large scaffold of quasi-random stable fixed point address states. Thus, we may view MESH and Vector-HaSH as fitting the diagram of Fig. 1 (right), except that the hidden layer in the diagram is replaced by a bipartite scaffold circuit (modular attractor network bidirectionally coupled to the densely connected layer) that computes nonlinear NN search to compute $\sigma(S({\bf K},{\bf Q}))$. In both cases, the memory values are reconstructed at the sensory input layer.
@@
The advantages of this construction are that (i) the address space is large, and (ii) the scaffold fixed points possess large and uniform basins of attraction. These properties allow many memories to be robustly addressed, with error correction and without collapse (blurring together the addresses of multiple memories). Such collapse is characteristic of the “memory cliff’ that afflicts many memory models (Amit et al.,, 1987; McCloskey and Cohen,, 1989): after a threshold number of inputs have been stored, retrieval performance crashes. In contrast, MESH and Vector-HaSH produce graceful degradation as multiple memories share overlapping addresses, much like human memory (e.g., Barnes and Underwood,, 1959). Remarkably, these models also outperform a flexible encoder trained to minimize reconstruction error (Radhakrishnan et al.,, 2020), suggesting that the entorhinal-hippocampal system may be a highly effective memory addressing system discovered by evolution.

### v2

@@ 1 Introduction
Despite the apparent fragility of memory, there is no decisive evidence that information, once stored, is ever permanently lost. The storage capacity of the brain of course must be finite, but that does not seem to be the principal limiting factor on memory performance. Rather, it is the retrieval process that fundamentally limits performance: the relevant information may be there, but cannot always be found 1, 2, 3, 4. Some of the evidence supporting this view will be summarized below.
A retrieval-oriented view of memory performance places the primary explanatory burden on how memories are addressed (i.e., how the retrieval system keeps track of storage locations), and how they are queried (i.e., how the retrieval system maps sensory cues to addresses). There is a long history of theorizing about these concepts in cognitive psychology and neuroscience 5, 6. Recently, the field of machine learning has developed its own analysis of these concepts, which form the basis of high-performing systems like transformers and fast weight programmers 7, 8. There is increasing recognition that a significant aspect of intelligence (in both natural and artificial systems) is effective information retrieval 9, 10, 11, 12, 13.
@@ 2.1 From correlations to kernels
One of the earliest formalizations of a key-value memory was Kohonen’s correlation matrix memory model 14, which was subsequently used by Pike 15 to explain a range of human memory phenomena. Here we slightly change the notation and terminology to bring this model into correspondence with more recent formalizations (Figure 1, left). Each input, indexed by $n$, consists of a key vector $\mathbf{k}_{n}$ and a value vector $\mathbf{v}_{n}$ (which we take to be row vectors). Intuitively, the keys encode information about memory addresses (hence we will refer to the set of possible keys as the address space), while the values encode information about memory content. These two representations are linked in memory by an “associator” matrix $\mathbf{M}$, which is initialized at 0 and incremented by the outer product of the key and value vectors after each input is presented:
@@
It is easy to see that this is just simple Hebbian learning between key and value units encoding the vector elements. In neurobiology, the standard interpretation is that $M_{ij}$ is encoded by the synaptic strength between neurons representing value element $j$ and key element $i$. The synaptic strength is increased when the two neurons fire coincidentally 16. Needless to say, learning in the brain is more complicated than this, but for present purposes we will take for granted the biological plausibility of Eq. 1 (see also 17 for a related, but more complex, learning rule).
@@
The correlation matrix memory is heteroassociative because $\mathbf{M}$ stores information about the relationship between two different kinds of objects or object properties (see also 18 for one of the earliest such models). If we impose the constraint that keys and values are the same, we get an autoassociative memory 19, 20, 21, 22. This idea is now most closely associated with the work of Hopfield 23, and such models (when the inputs are binary) are typically referred to as Hopfield networks.
@@
In neural network terms, this is equivalent to activating pattern $\mathbf{q}$ on the key units, which produces an activity pattern in downstream value units based on the learned synaptic strengths ($\mathbf{M}$). The retrieved value vector $\hat{\mathbf{v}}$ combines stored values associated with keys that are similar to the query vector. To see this, it is helpful to rewrite Eq. 2 in its “dual” form 24, 25:
@@
where $\{\alpha_{n}\}$ is a set of “attention weights” computed by $\alpha=\sigma(S(\mathbf{K},\mathbf{q}))=\mathbf{q}\mathbf{K}^{\top}$, and $\mathbf{K}$ is the matrix consisting of all $N$ key vectors (i.e., its rows are $\mathbf{k}_{n}$). The function $S(\cdot,\cdot)$ is a similarity kernel (linear in this case), expressing the match between keys and queries, and $\sigma(\cdot)$ is a separation operator (the identity function in this case), pushing apart similar memories. The retrieved value vector is thus a weighted combination of stored values, where the attention weights correspond to the similarity between the query and each key, mapped through a separation operator.
The dual form is useful because it shows how the model can be straightforwardly generalized by considering other similarity kernels and separation operators. Any positive semidefinite kernel can be expressed as the inner product in some vector space 26, which means that alternative similarity functions can be constructed by mapping the keys and queries through a feature transform $\phi(\cdot)$ applied to each row:
@@
It is also possible to construct kernels with infinite-dimensional vector spaces by working directly with an explicitly defined similarity kernel, such as the widely used radial basis function kernel, which Tsai and colleagues 27 found to produce the best results.
The generalized correlation matrix memory can implement the celebrated self-attention mechanism in transformers 7:
@@
where $D$ is the dimensionality of the key/value vectors, and the softmax function is used as the separation operator. Here we have used $\tilde{\alpha}$ to denote the attention weights prior to the application of the separation operator. When $\sigma(\cdot)$ is set to the identity function, we recover linearized attention 25, which can be expressed as a special form of recurrent neural network known as a fast weight programmer 8, 28. Linearized attention has a potential computational advantage over softmax attention, because its recurrent form allows for linear-complexity sequence processing, while giving up the parallel computation property of the attention form (whose complexity, however, is quadratic in the number of sequence elements). Several studies 29, 30 have shown how to rigorously approximate softmax attention using random features.
As shown by Millidge and colleagues 31, a variety of classical memory models can also be derived by different choices of similarity kernel and separation function. For example, sparse distributed memory 32 can be obtained by setting $\sigma(\cdot)$ to a threshold function, and dense associative memory 33 can be obtained by setting $\sigma(\cdot)$ to a rectified polynomial.
In the noiseless setting, the ideal separation function is the max operator, since this will always return the stored value associated with the matching key 31. However, this separation function is not robust to noise—small perturbations can cause large retrieval errors. Thus, the design of a memory system needs to balance separability and robustness. In addition, the memory may be used for generalization to new inputs 7, 34, 35, in which case perfect matching is not the goal. For example, key-value memories have been used extensively for question-answering tasks 36, 37, where $\mathbf{q}$ represents a question and $\mathbf{v}$ represents an answer. The system stores a set of question-answer pairs and tries to use this knowledge base to answer novel questions.
@@ 2.2 Representational structure
In some systems, the key and query mappings are assumed to be fixed, either randomly or based on some regular structure. In these systems, the address space is conceived as a kind of “scaffold” for the indexing of information content. In the sparse distributed memory model 32, the scaffold is random, so that similar keys do not systematically index similar values. This mimics (approximately) the organization of random access memory in digital computers. A familiar example of a structured scaffold is the alphabetical index found at the end of books. In the Hopfield network 23 and related autoassociative memory models, the scaffold is identical to the value space. This property endows the models with content addressability 38: memory addresses are accessed by matching values directly to queries. Most psychological and neural models of memory share this property, though they implement it in different ways.
Some models assume that the scaffold is randomly constructed in such a way that it carries a structural imprint. For example, if inputs are assigned to random addresses, but these addresses change slowly over time, then inputs experienced nearby in time will encoded as more “similar” in the address space. Temporal autocorrelation, in the absence of additional structure, is able to account for many aspects of human 39 and animal 40 memory.
@@ 2.3 The ubiquity of key-value memory
The work reviewed above focused on explicit constructions of key-value memory. It turns out that other models are sometimes implicitly equivalent. Irie and colleagues 12 showed that linear layers trained by gradient descent (ubiquitous in many machine learning models) can also be expressed as key-value memories. Here we briefly review this theorem.
@@
where $\mathbf{W}_{0}$ is the initial weight matrix, and $\mathbf{e}_{n}=-\eta_{n}(\nabla_{\mathbf{y}}\mathcal{L})_{n}$ is the error signal with learning rate $\eta_{n}$ and loss function $\mathcal{L}$. Generalizing a classic result on kernel machines 41, Irie and colleagues 12 showed that this construction is equivalent (in the sense of producing the same output for a given input) to the following linear key-value memory:
@@
where (using our earlier notation) $\alpha=\mathbf{q}\mathbf{K}^{\top}$ with $\mathbf{v}_{n}=\mathbf{e}_{n}$, $\mathbf{k}_{n}=\mathbf{x}_{n}$, and $\mathbf{q}=\mathbf{x}$, given an arbitrary input $\mathbf{x}$. Thus, linear layers effectively memorize their experienced error patterns, computing their outputs as linear functions of these memories. This interpretation is intriguing in light of the fact that errors are particularly salient in human memory 42, 43, 44, 45, 46.
The linear layers retain information about all training inputs—they never “forget” 12. However, retrieval access may be (transiently) lost. We discuss evidence for this idea from psychology and neuroscience below.
@@ 3 Neurobiological substrates
While key-value memories are loosely inspired by the brain, it remains an open question how to implement them in a biologically plausible manner. Eq. 1 is a (somewhat) biologically plausible rule for learning associations between keys and values, but we still need rules for storing the keys and values themselves. In the section on representation learning, we described the widely used approach of modeling keys, queries, and values as linear mappings from input vectors (which themselves could be learned). This implies additional learning rules for the linear mappings. One could posit that they are learned by backpropagating the errors from whatever objective function is being optimized for a downstream task 17. This would require a biologically plausible approximation of the backpropagation algorithm 47.
Kozachkov and colleagues 48 have proposed an architecture based on the “tripartite synapse” consisting of pre-synaptic and post-synaptic neurons modulated by an astrocyte (a type of glial cell). In particular, they posited that this motif applies to the hidden-to-output synapses in the three-layer network implementation of key-value memory. The activation of each astrocyte process is modeled as a linear function of the hidden unit activations. They showed how the astrocyte processes collectively compute the similarity function $S(\mathbf{K},\mathbf{q})$, which then multiplicatively modulate the hidden-to-output weights so that the network as a whole implements the transformer self-attention described earlier.
Alternatively, Tyulmankov and colleagues 49 have proposed a non-Hebbian learning rule for key learning. They view the key-value memory as a three-layer neural network, where the input $\mathbf{x}$ (first layer) is transformed into a pattern of attention $\boldsymbol{\alpha}$ (hidden layer), which is finally transformed into retrieved values $\hat{\mathbf{v}}$ (output layer); see the right panel of Figure 1. Each hidden layer unit represents a “slot” to which a single input (or possibly multiple inputs) gets assigned. In this view, the key matrix $\mathbf{K}$ corresponds to the input-hidden synaptic strengths, and the value matrix $\mathbf{V}$ (where row $n$ corresponds to value vector $\mathbf{v}_{n}$) corresponds to the hidden-output synaptic strengths. The proposed learning rule for the synapse connecting input unit $j$ to hidden unit $i$ is (simplifying slightly, and dropping time indices):
@@
where $\mu\in\{0,1\}$ is a global third factor (possibly a neuromodulator like acetylcholine or dopamine) and $\gamma_{i}\in\{0,1\}$ is a local third factor (possibly a dendritic spike). The factors are binary to promote sparsity in the hidden layer. The local third factor tags the least-recently-used hidden unit as eligible for plasticity. The authors point out that the key learning rule resembles behavioral time scale plasticity 50, which has been observed in the hippocampus. The possibility that key learning occurs in the hippocampus will be considered in detail below.
For value learning, Tzyulmankov and colleagues clamp the output layer to the target value $\mathbf{v}$ and then update the synapse connecting hidden unit $i$ to output unit $m$ according to:
@@
Using a similar architecture (but with the important addition of recurrence), Whittington and colleagues 51 have shown that the Tolman-Eichenbaum Machine 52, a model of the entorhinal-hippocampal system, implements a form of key-value memory. The hippocampus, according to this model, stores conjunctive representations of sensory inputs from lateral entorhinal cortex and a “structural code” provided by cells in medial entorhinal cortex, which are updated recurrently. Conjunctive representations in the hippocampus function as both the keys and the values (i.e., an autoassociative memory). In a 2D open field, the model can reproduce hippocampal “place cells” (firing fields localized to a particular spatial location), while the structural code generates grid cell-like responses (grid cells fire with hexagonal periodicity as an animal moves through space 53). In environments of different dimensions, topologies, and geometries, the structural code need not be grid-like. Whittington and colleagues note that the structural code in their model functions as a form of “position encoding”—widely used in transformers and other sequence models—where the encodings (i.e., addresses) are adapted to the structure of space. This is a critical aspect of the model: the structure-sensitive properties of their modeled hippocampal and entorhinal cells would not have been obtained using the fixed positional encodings typically used in regular transformers. Related ideas have proven successful in the machine learning literature 54.
Box 1: Memory with error-correcting key-query matching
@@
Vector-HaSH 55 and MESH 56 are tripartite networks consisting of an input (sensory) layer $s$, a densely connected layer $d$, and a layer consisting of a modular set of fixed recurrent attractor networks $a$. These models dissociate content storage from query-matching and directly translate to a key-query-value formulation. Define the tripartite network weights to be ${\bf W}_{sd},{\bf W}_{da},{\bf W}_{ad},{\bf W}_{ds}$, and the attractor, dense layer, and sensory states to be $\{{\bf a}\},\{{\bf d}\},\{{\bf s}\}$, respectively. Given some sensory input ${\bf s}$, the network retrieves the nearest relevant stored state $\hat{\bf s}$ by performing the following operations (treating the dense layer responses as linear, though in practice they are nonlinear): $\hat{{\bf s}}=\phi({\bf s}{\bf W}_{sd}{\bf W}_{da},\{{\bf a}\}){\bf W}_{ad}{\bf W}_{ds}.$ Here $\phi(.,.)$ represents an efficient nearest-neighbor (NN) operation performed by the scaffold (combined dense and attractor layers), to select the closest attractor state for the given input. The readout weights $W_{ad},W_{ds}$ reconstruct the sensory memory from the selected attractor state. This model maps to the key-value system of Figure 1B: the output of the attractor layer is a (denoised) query ${\bf q}_{n}=\phi({\bf s}{\bf W}_{sd}{\bf W}_{da},\{{\bf a}\})\equiv{\bf x}_{n}$ and the key and value matrices are given by the attractor-to-dense and dense-to-sensory weights, repsectively, $\mathbf{K}={\bf W}_{ad}^{\top},\ ~{}\mathbf{V}={\bf W}_{ds}.$ Thus, the scaffold computes a similarity $\sigma(S({\bf K},{\bf q}))$ and the network produces the output $\hat{{{\bf v}}}_{n}=\hat{{\bf s}}_{n}=\alpha_{n}{\bf W}_{ds}=\phi({\bf s}_{n}{\bf W}_{sd}{\bf W}_{da},\{{\bf a}\}){\bf W}_{ad}{\bf W}_{ds}=\sigma(S({\bf K},{\bf q}_{n})){\bf V}$ where the attention weights (Fig. 1B’s hidden states) are $\alpha_{n}={\bf x}_{n}{\bf W}_{ad}$. The value weights (from the dense to the sensory layer) and part of the query weights (from the sensory to the dense layer) are trained by a Hebbian rule, so that ${\bf W}_{sd}=\sum_{n}{\bf s}_{n}^{\top}{\bf d}_{n}$ and ${\bf W}_{ds}=\sum_{n}{\bf d}_{n}^{\top}{\bf s}_{n}$ the weights ${\bf W}_{da}$ are random and fixed, and ${\bf W}_{ad}$ are also held fixed (after being set as a pseudoinverse on the random ${\bf W}_{da}$, so these may also be viewed as random). Thus, the query weights are partly random, fixed, and coupled with the keys, and partly learned through Hebbian plasticity. The choice of modular attractor states and random projections (key weights) from attractor to dense layer pre-defines both well-separated keys and a NN operation for robust error-correction and matching of noisy queries with keys. Though learning is heteroassociative and values are stored separately from keys and queries, unlike Hopfield networks, the networks are content-addressable.
@@
In contrast to the adaptive addresses used in the Tolman-Eichenbaum Machine, Sharma, Chandra and colleagues 56, 55 proposed neurobiologically motivated models (MESH and Vector-HaSH) that use random projections of modular fixed point (attractor) networks to address memories (see Box 3 for details). The fixed points within each modular attractor, which are fixed and not learned, may be sparse and random 56 or possess a 2-dimensional geometry 55 such as the hexagonally periodic activity patterns of grid cells. Consisent with the non-learned assumption, grid cells exhibit a fixed relational organization regardless of environment and behavioral state 57, 58, 59, 60. The same relationships are re-used to encode the organization of non-spatial variables 61, 62, 63. The fixed modular attractor networks in the models randomly and densely project into a larger network (the hippocampus in Vector-HaSH) with self-consistent return projections, forming a large scaffold of quasi-random stable fixed point address states. Thus, we may view MESH and Vector-HaSH as fitting the diagram of Fig. 1 (right), except that the hidden layer in the diagram is replaced by a bipartite scaffold circuit (modular attractor network bidirectionally coupled to the densely connected layer) that computes nonlinear robust NN search to compute $\sigma(S({\bf K},{\bf Q}))$, with keys that are random and fixed. In Vector-HaSH, keys reside in hippocampus and the memory values are reconstructed at the sensory input layer (cortex).
@@
The advantages of this construction, beyond standard key-value constructions, are that (i) the address space is large, and (ii) the scaffold fixed points possess large and uniform basins of attraction. These properties allow many memories to be robustly addressed, with error correction and without the “memory cliff’ that afflicts many memory models combine keys and values 64, 65 or learn the keys 55: after a threshold number of inputs have been stored, retrieval performance crashes. In contrast, MESH and Vector-HaSH produce graceful degradation as multiple memories share overlapping addresses, much like human memory 66. Remarkably, these models also outperform a flexible encoder trained to minimize reconstruction error 67, suggesting that the entorhinal-hippocampal system may be a highly effective memory addressing system discovered by evolution.